Find three consecutive even integers if the product of the second and third is eighty eight more than four times the first.
So I have (x+2)(x+4)= 4(x)+ 84. Then when I combine like terms I have x^2+2x-76. Is this correct? If so, how does it factor?

Answer by rajagopalan(174)

Solution
you have (x+2)(x+4)= 4(x)+ 88.
x^2+6x+8=4x+88 (not eighty four as u assumed)
x^2+2x-80=0
x^2-8x+10x-80=0
x(x-8)+10(x-8)=0
(x+10)(x-8)=0
gives x=-10 or 8
take 8 as positive answer
0thers are 10 and 12 obviously
good day
cheers